# Layered, Overlapping, and Inconsistent: A Large-Scale Analysis of the Multiple Privacy Policies and Controls of U.S. Banks

### v1

@@ Abstract.
Privacy policies are often complex. An exception is the two-page standardized notice that U.S. financial institutions must provide under the Gramm-Leach-Bliley Act (GLBA). However, banks now operate websites, mobile apps, and other services that involve complex data sharing practices that require additional privacy notices and do-not-sell opt-outs. We conducted a large-scale analysis of how U.S. banks implement privacy policies and controls in response to GLBA; other federal privacy policy requirements; and the California Consumer Privacy Act (CCPA), a key example for U.S. state privacy laws. We focused on the disclosure and control of a set of especially privacy-invasive practices: third-party data sharing for marketing-related purposes. We collected privacy policies for the 2,067 largest U.S. banks, 45.3% of which provided multiple policies. Across disclosures and controls within the same bank, we identified frequent, concerning inconsistencies—such as banks indicating in GLBA notices that they do not share with third parties but disclosing sharing elsewhere, or using third-party marketing/advertising cookies without disclosure. This multiplicity of policies, with the inconsistencies it causes, may create consumer confusion and undermine the transparency goals of the very laws that require them. Our findings call into question whether current policy requirements, such as the GLBA notice, are achieving their intended goals in today’s online banking landscape. We discuss potential avenues for reforming and harmonizing privacy policies and control requirements across federal and state laws.
Privacy, finance, privacy notice, opt-out, third-party sharing.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 13–17, 2025; Taipei, Taiwanisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Human and societal aspects of security and privacyccs: Social and professional topics Privacy policies
@@ 1. Introduction
In the U.S., consumer privacy laws rely on the “notice and choice” framework (Federal Trade Commission, 2008), which requires businesses to provide privacy notices and opt-out choices, so that consumers can understand data practices and exercise privacy choices. Numerous studies have shown that this framework neither effectively informs consumers nor supports their ability to express choices (e.g., Cranor, 2012a, 2024; Cate, 2016). Design efforts have sought to make privacy information more accessible and comprehensible (Schaub et al., 2015; Kelley et al., 2009; Zhang et al., 2022). A notable example is the standardized short-form notice (see Figure 1) that financial institutions must provide under the federal Gramm-Leach-Bliley Act (GLBA) (Federal Trade Commission, 2002; Garrison et al., 2012).
@@
Given the range of privacy regulations and the differences in their scope, definitions, and requirements, U.S. banks have started providing multiple notices and opt-out choices in addition to the GLBA short-form notice. We conducted a large-scale analysis of the privacy policies and respective privacy controls for the 2,073 largest commercial banks in the U.S., which collectively hold 97.3% of all assets of FDIC-insured commercial banks. We examined whether a given bank provides multiple privacy policies (GLBA, general, mobile, cookie, and CCPA policies) and controls (GLBA, cookie, and CCPA opt-outs), and whether inconsistencies exist across these policies that could mislead or confuse consumers. We focused specifically on third-party data sharing for marketing or advertising and related opt-outs because people often find these practices concerning (Rao et al., 2016; Weinshel et al., 2019; Al-Ameen et al., 2020), as they constitute violations of contextual integrity  (Nissenbaum, 2011). Our research questions are:
@@
    How do banks provide privacy opt-outs regarding third-party sharing for marketing purposes that are required by different regulations?
Summary of findings. We found privacy policies for 2,069 banks, 45.3% of which provided multiple privacy policies, most commonly a GLBA notice in combination with a general or mobile privacy policy. The combined privacy policy content for each bank requires a median reading level equivalent to college education, with larger banks providing lengthier and more difficult-to-read text. We found concerning inconsistencies: 55.2% of banks with multiple privacy policies indicated in their GBLA notice that they would not share data with third parties for marketing purposes, yet disclosed such sharing in their other policies. Interestingly, in fewer cases, we found the opposite: sharing was indicated in the GLBA notice while the bank indicated in other disclosures (often CCPA) not to sell/share data. Many banks that disclosed data-sharing practices did not offer corresponding required opt-outs.
The multiplicity of policies and identified inconsistencies shows that for many banks, the GLBA notice no longer provides a full representation of their third-party sharing practices. Consumers now must navigate multiple privacy documents, with the documents’ varying scopes in mind, to learn about and manage their data and understand their privacy choices. Our findings highlight that the narrowly-scoped GLBA notice may mislead consumers, and that the layering of different disclosure requirements can undermine the transparency goals of the very laws that require them. We discuss opportunities for regulatory reform that could reduce duplication, resolve inconsistencies across notices, and ultimately make privacy information more accessible and actionable for consumers.
@@ 2. Background
We discuss privacy notice and opt-out requirements of GLBA, other privacy notice requirements, and GLBA.
@@ 2.1. The Gramm-Leach-Bliley Act
The Gramm-Leach-Bliley Act of 1999 (GLBA) requires financial institutions to disclose their information collection and sharing practices annually as well as inform customers of their right to opt out of certain sharing practices (Congress, 1999). The GLBA narrowly covers “nonpublic personal information” related to providing financial products or services (Federal Trade Commission, 2002; United States Congress, 2010), such as a consumer’s name, income, and social security number.
A two-page GLBA model privacy form (Securities and Commission, 2009) prescribes a standardized layout with pink-bracketed text to be customized for an institution’s data practices (see Figure 1). Although GLBA notices were historically delivered by postal mail, many banks now send these notices via email and provide them on their websites in PDF format.
Figure 1. 2-page GLBA model privacy form (Securities and Commission, 2009). We analyzed the (a) data-sharing table, (b) opt-out box, (c) OII box. Blue annotations were added for clarity, not part of the template.
The model privacy form highlights how bank customers’ financial information is collected, shared, and protected in a table-based format. It was consumer tested for judgment quality, perceptual accuracy, and reading ease (Group, 2006; Garrison et al., 2012; Levy and Hastak, 2008). On the first page, the “Why?,” ”What?,” and ”How?” text boxes summarize disclosure requirements as well as five examples of personal information types, chosen from a pre-defined list (Consumer Financial Protection Bureau, [n. d.]), that the institution collects and shares. Our analysis focuses on the data-sharing table (data-sharing table) (Figure 1(a)), which lists at most seven pre-defined data-sharing practices distinguished by sharing purposes and informs consumers about whether they can limit each sharing. Among the seven purposes, the “For our affiliates to market to you” purpose may be omitted in certain cases, while the others are required to be listed.
An institution must provide an opt-out option for the last three of these purposes. Respective opt-out instructions are described in a “To limit our sharing” box on the first page (opt-out box, Figure 1(b)). We found that many financial institutions use the “Other important information” box (OII box) on the second page to point to additional data practice disclosures and opt-out rights under state-level privacy laws (Figure 1(c)). The disclosures refer to affiliates, which are financial or nonfinancial companies under common ownership or control with the disclosing institution; nonaffiliates, which are unrelated third parties; and joint marketers, which can be affiliates and nonaffiliates (Consumer Financial Protection Bureau, [n. d.]).
@@ 2.3. California Privacy Laws
California has a state privacy law for the financial industry, the California Financial Information Privacy Act (CalFIPA) (California Legislature, 2003), which requires explicit consent for certain types of marketing-related third-party sharing. In addition, the California Consumer Privacy Act (CCPA) is a comprehensive privacy law encompassing almost all industries. Businesses are subject to CCPA if they operate in California; collect, sell, or share consumer personal information; and meet certain thresholds (California Privacy Protection Agency, 2023; Information, [n. d.]). CCPA defines “personal information” more broadly than GLBA’s “nonpublic personal information,” covering any data that identifies, relates to, or can reasonably be linked to a consumer or their household. This includes consumers’ names, social security numbers, email addresses, records of purchased products, browsing history, location data, etc. Under CCPA, businesses that sell or share personal information with third parties are required to provide a notice of the right to opt-out of the sale/sharing and a method for doing so (California Legislature, 2025). Sharing under CCPA specifically refers to disclosures made for “cross-context behavioral advertising” with third parties, excluding service providers and contractors (California Legislature, 2025). Compared to GLBA, CCPA specifies required content in detail but does not mandate notice formats. We observed that some banks included the notice of the right to opt out of sale/sharing within their general privacy policy, while others used a standalone CCPA policy, both with varying formats.
@@
The CCPA includes a carve-out for data covered under GLBA, meaning that personal information collected, processed, sold, or disclosed pursuant to the GLBA (i.e., as part of financial transactions) is exempt from CCPA requirements. However, banks are subject to CCPA for other personal information they process. For example, website tracking data, such as cookies that monitor browsing behavior for retargeting purposes, is outside of GLBA but covered under CCPA. Similarly, a bank website’s use of third-party tracking scripts for cross-context behavioral advertising falls under CCPA. Partial overlaps and gaps like these create a patchwork in which different information of the same consumer held by a bank is governed by multiple regulations, with associated data practices disclosed in different privacy policies.
@@ 3. Related Work
We present related work on inconsistencies in privacy disclosures and practices, readability of privacy policies, usability of privacy controls, and legal compliance analysis.
@@
Readability of Privacy Policies. Extensive research has examined the readability of privacy policies, identifying persistent clarity issues despite regulatory efforts, and finding that regulation has had a mixed effect on privacy policy transparency (Birrell et al., 2023). For example, Chen et al.’s (Chen et al., 2021) analysis of the privacy policies of 95 popular websites found that despite the CCPA’s mandate for clear privacy disclosures, privacy policies varied significantly in both the level of detail provided and in how key CCPA definitions like “sale”, “valuable consideration”, or “business purpose” are interpreted by businesses. Their survey found that many consumers found it difficult to fully understand how their data is collected and shared. Similar concerns arise in studies examining the impact of the GDPR on privacy disclosures. Kretschmer et al. (Kretschmer et al., 2021) and Degeling et al. (Degeling et al., 2018) found that while transparency has improved since the introduction of the GDPR, usability challenges remain, especially in complex interface designs that limit user agency. Wagner et al. (Wagner, 2023a) also found that privacy policies remain difficult to read, with legal reforms prompting only incremental improvements in readability. These findings suggest that despite regulatory intentions for transparency, privacy policies often remain too complex, inconsistent, or difficult to navigate, which limits their effectiveness in truly informing consumers (Cranor, 2012b). Rather than examining the full scope of privacy policies, we focus on consumers’ information needs related to exercising privacy controls—which typically pertain to limiting data sharing—and analyze how data-sharing practices are disclosed across privacy policies of the same bank.
@@
Compliance with Privacy Regulations. Businesses’ privacy policies and data practices often fail to meet regulatory requirements (Van Nortwick and Wilson, 2022; Aziz and Wilson, 2024; Tran et al., 2024a, b; Samarin et al., 2023). Cranor et al.’s (Cranor et al., 2016) large-scale study of the GLBA notices from financial institutions found that many failed to provide required opt-out mechanisms, and some incorrectly stated that consumers could not limit certain types of data sharing that the GLBA permits them to restrict. Our analysis of GLBA notices ten years later than Cranor et al. (Cranor et al., 2016) found lower percentages of non-compliance. Similarly, studies examining CCPA and GDPR compliance (Tran et al., 2024a; O’Connor et al., 2021; Samarin et al., 2023) reveal that businesses frequently fail to implement legally required opt-out links and cookie controls. Even when these mechanisms are provided, manipulative design tactics are often used to discourage consumers from exercising their rights, in direct violation of CCPA’s restrictions (Tran et al., 2024b). Additionally, Aziz et al. (Aziz and Wilson, 2024) and Zimmeck et al. (Zimmeck et al., 2023) showed that automated privacy signals like Global Privacy Control (GPC), which are intended to provide consumers with an automated way to opt out, are frequently ignored by websites, despite legal mandates to honor them. Similarly, we found that some banks in our set ignored GPC signals even though they disclosed sale/sharing under CCPA definition. Beyond interface-level manipulations, compliance gaps persist in actual data practices. Studies by Matte et al. (Matte et al., 2020) and Zhou et al. (Zhou et al., 2023) found that many apps and websites continue to collect personal data without proper disclosure, often using pre-selected consent options that subtly steer users toward agreement. The growing complexity of sector-specific and state-level privacy laws makes it increasingly difficult to maintain compliance and coherence. Our study examines the interplay between different compliance requirements through a consistency analysis, as consistency, clarity, and accessibility both within and across privacy documents provided by a single organization shape a consumer’s understanding and control of the use of their data.

### v2

@@ Abstract.
Privacy policies are often complex. An exception is the two-page standardized notice that U.S. financial institutions must provide under the Gramm-Leach-Bliley Act (GLBA). However, banks now operate websites, mobile apps, and other services that involve complex data sharing practices that require additional privacy notices and do-not-sell opt-outs. We conducted a large-scale analysis of how U.S. banks implement privacy policies and controls in response to GLBA; other federal privacy policy requirements; and the California Consumer Privacy Act (CCPA), a key example for U.S. state privacy laws. We focused on the disclosure and control of a set of especially privacy-invasive practices: third-party data sharing for marketing-related purposes. We collected privacy policies for the 2,067 largest U.S. banks, 45.2% of which provided multiple policies. Across disclosures and controls for the same bank, we identified frequent, concerning inconsistencies—53.8% of banks with multiple privacy policies indicated in GLBA notices that they do not share with third parties but disclosed sharing in other policies. This multiplicity of policies, with the inconsistencies it causes, may create consumer confusion and undermine the transparency goals of the very laws that require them. Our findings call into question whether current policy requirements, such as the GLBA notice, are achieving their intended goals in today’s online banking landscape. We discuss potential avenues for reforming and harmonizing privacy policies and control requirements across federal and state laws.
Privacy, finance, privacy notice, privacy opt-out, third-party sharing.
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765072isbn: 979-8-4007-1525-9/2025/10ccs: Security and privacy Human and societal aspects of security and privacyccs: Social and professional topics Privacy policies
@@ 1. Introduction
In the U.S., consumer privacy laws rely on the “notice and choice” framework (Federal Trade Commission, 2008), which requires businesses to provide privacy notices and opt-out choices, so that consumers can understand data practices and exercise privacy choices. Numerous studies have shown that this framework neither effectively informs consumers nor supports their ability to express choices (e.g., Cranor, 2012a, 2024; Cate, 2016). Design efforts have sought to make privacy information more accessible and comprehensible (Schaub et al., 2015; Kelley et al., 2009; Zhang et al., 2022). A notable example is the standardized short-form notice (see Figure 1) that U.S. financial institutions must provide under the federal Gramm-Leach-Bliley Act (GLBA) (Federal Trade Commission, 2002; Garrison et al., 2012).
@@
Given the range of privacy regulations and the differences in their scope, definitions, and requirements, U.S. banks have started providing multiple notices and opt-out choices in addition to the GLBA short-form notice. We conducted a large-scale analysis of the privacy policies and respective privacy controls for the 2,073 largest commercial banks in the U.S., which collectively hold 97.3% of all assets of FDIC-insured commercial banks. We examined whether a given bank provides multiple privacy policies (GLBA, general, mobile, cookie, and CCPA policies) and controls (GLBA, cookie, and CCPA opt-outs), and whether inconsistencies exist across these policies that could mislead or confuse consumers. We focused specifically on third-party data sharing for marketing, which we considered to include both marketing/advertising and analytics/research purposes more broadly, along with the related opt-outs, because people often find these practices concerning (Rao et al., 2016; Weinshel et al., 2019; Al-Ameen et al., 2020), as they constitute violations of contextual integrity (Nissenbaum, 2011). Our research questions were:
@@
    How do banks provide privacy opt-outs regarding third-party sharing for marketing purposes, as required by different regulations?
Summary of findings. We found privacy policies for 2,069 banks, 44.2% of which provided multiple privacy policies, most commonly a GLBA notice in combination with a general or mobile privacy policy. The combined privacy policy content for each bank requires a median reading level equivalent to a college education, with larger banks providing lengthier and more difficult-to-read text. We identified two distinct types of inconsistency across a bank’s privacy disclosures. Concerningly, many banks appear more privacy protective in the GLBA notice than in other disclosures: 53.8% of banks with multiple privacy policies indicated in their GLBA notice that they would not share data with third parties for marketing purposes, yet disclosed such sharing in their other policies. The opposite inconsistency type was less concerning, though still notable: 36.8% of banks with multiple privacy policies indicated sharing in the GLBA notice, while in CCPA-related disclosures they stated that they do not sell/share data, possibly reflecting CCPA’s stricter limits on data sharing. Many banks that disclosed data-sharing practices did not offer the required opt-outs, and most banks had third-party marketing cookies on their websites without disclosing them or offering opt-out controls.
The multiplicity of policies and identified inconsistencies shows that for many banks the short-form GLBA notice no longer provides a full representation of their third-party sharing practices. Consumers must now navigate multiple privacy documents, with the documents’ varying scopes in mind, to learn about and manage their data and understand their privacy choices. Our findings highlight that the narrowly-scoped GLBA notice may mislead consumers, and that the layering of different disclosure requirements can undermine the transparency goals of the very laws that require them. We discuss concrete opportunities for regulatory reform that could reduce duplication, resolve inconsistencies across notices, and ultimately make privacy information more accessible and actionable for consumers.
@@ 2. Background
We discuss privacy notice and opt-out requirements of GLBA, other privacy notice requirements, and California privacy laws.
@@ 2.1. The Gramm-Leach-Bliley Act
The Gramm-Leach-Bliley Act of 1999 (GLBA) requires financial institutions to disclose their information collection and sharing practices annually as well as inform customers of their right to opt out of certain sharing practices (Congress, 1999). The GLBA narrowly covers “nonpublic personal information” related to providing financial products or services (Federal Trade Commission, 2002; United States Congress, 2010), such as a consumer’s name, income, and social security number. A two-page GLBA model privacy form (Securities and Commission, 2009) prescribes a standardized layout with pink-bracketed text to be customized for an institution’s data practices (see Figure 1). Although GLBA notices were historically delivered by postal mail, many banks now send these notices via email and provide them on their websites in PDF format.
Figure 1. 2-page GLBA model privacy form (Securities and Commission, 2009). We analyzed the (a) data-sharing table, (b) opt-out box, (c) OII box (blue annotations added for clarity, not part of the template).
The model privacy form highlights how bank customers’ financial information is collected, shared, and protected in a table-based format. It was consumer tested for judgment quality, perceptual accuracy, and reading ease (Group, 2006; Garrison et al., 2012; Levy and Hastak, 2008). On the first page, the “Why?,” “What?,” and “How?” text boxes summarize disclosure requirements as well as five examples of personal information types, chosen from a pre-defined list (Consumer Financial Protection Bureau, [n. d.]), that the institution collects and shares. Our analysis focuses on the data-sharing table (data-sharing table; see Figure 1(a)), which lists at most seven pre-defined data-sharing practices distinguished by sharing purposes and informs consumers about whether they can limit each sharing. Among the seven purposes, the “For our affiliates to market to you” purpose may be omitted in certain cases, while the others are required to be listed.
An institution must provide an opt-out option for the last three of these purposes. Respective opt-out instructions are described in a “To limit our sharing” box on the first page (opt-out box, see Figure 1(b)). We found that many financial institutions use the “Other important information” box (OII box, see Figure 1(c)) on the second page to point to additional data practice disclosures and opt-out rights under state-level privacy laws. The disclosures refer to affiliates, which are financial or nonfinancial companies under common ownership or control with the disclosing institution; nonaffiliates, which are unrelated third parties; and joint marketers, which can be affiliates and nonaffiliates (Consumer Financial Protection Bureau, [n. d.]).
@@ 2.3. California Privacy Laws
California has a state privacy law for the financial industry, the California Financial Information Privacy Act (CalFIPA) (California Legislature, 2003), which requires explicit consent for certain types of marketing-related third-party sharing. In addition, the California Consumer Privacy Act (CCPA) is a comprehensive privacy law encompassing almost all industries. Businesses are subject to CCPA if they operate in California; collect, sell, or share consumer personal information; and meet certain thresholds (California Privacy Protection Agency, 2023; Information, [n. d.]). CCPA defines “personal information” more broadly than GLBA’s “nonpublic personal information,” covering any data that identifies, relates to, or can reasonably be linked to a consumer or their household. This includes consumers’ names, social security numbers, email addresses, records of purchased products, browsing history, location data, etc.
Under CCPA, businesses that sell or share personal information with third parties are required to provide a notice of the right to opt-out of the sale/sharing and a method for doing so (California Legislature, 2025). Sharing under CCPA specifically refers to disclosures made for “cross-context behavioral advertising” with third parties, excluding service providers and contractors (California Legislature, 2025). Compared to GLBA, CCPA specifies required content in detail but does not mandate notice formats. We observed that some banks included the notice of the right to opt out of sale/sharing within their general privacy policy, while others used a standalone CCPA policy, both with varying formats.
@@
The CCPA includes a carve-out for data covered under GLBA, meaning that personal information collected, processed, sold, or disclosed pursuant to the GLBA (i.e., as part of financial transactions) is exempt from CCPA requirements. However, banks are subject to CCPA for other personal information they process. For example, website tracking data, such as cookies that monitor browsing behavior for retargeting purposes, is outside of GLBA’s scope but covered under CCPA. Similarly, a bank website’s use of third-party tracking scripts for cross-context behavioral advertising falls under CCPA. Partial overlaps and gaps like these create a patchwork in which different information of the same consumer held by a bank is governed by multiple regulations, with associated data practices disclosed in different privacy policies.
@@ 3. Related Work
We discuss related work on inconsistencies in privacy disclosures and practices, readability of privacy policies, usability of privacy controls, and legal compliance analysis.
@@
Readability of privacy policies. Extensive research has examined the readability of privacy policies, identifying persistent clarity issues despite regulatory efforts, and finding that regulation has had a mixed effect on privacy policy transparency (Birrell et al., 2023). For example, Chen et al.’s (Chen et al., 2021) analysis of the privacy policies of 95 popular websites found that despite the CCPA’s mandate for clear privacy disclosures, privacy policies varied significantly in both the level of detail provided and in how key CCPA definitions like “sale,” “valuable consideration,” or “business purpose” are interpreted by businesses. Their survey found that many consumers found it difficult to fully understand how their data is collected and shared. Similar concerns arise in studies examining the impact of the GDPR on privacy disclosures. Kretschmer et al. (Kretschmer et al., 2021) and Degeling et al. (Degeling et al., 2018) found that while transparency has improved since the introduction of the GDPR, usability challenges remain, especially in complex interface designs that limit user agency. Wagner et al. (Wagner, 2023a) also found that privacy policies remain difficult to read, with legal reforms prompting only incremental improvements in readability. These findings suggest that despite regulatory intentions for transparency, privacy policies often remain too complex, inconsistent, or difficult to navigate, which limits their effectiveness in truly informing consumers (Cranor, 2012b). Rather than examining the full scope of privacy policies, we focus on consumers’ information needs related to exercising privacy controls—which typically pertain to limiting data sharing—and analyze how data-sharing practices are disclosed across privacy policies of the same bank.
@@
Compliance with privacy regulations. Businesses’ privacy policies and data practices often fail to meet regulatory requirements (Van Nortwick and Wilson, 2022; Aziz and Wilson, 2024; Tran et al., 2024a, b; Samarin et al., 2023). Cranor et al.’s (Cranor et al., 2016) large-scale analysis of financial institutions’ GLBA notices found that, in 2014, many failed to provide required opt-out mechanisms, and some incorrectly stated that consumers could not limit certain types of data sharing that the GLBA permits them to restrict. Our analysis of GLBA notices found lower percentages of non-compliance than Cranor et al. reported a decade ago (Cranor et al., 2016). We further expand on Cranor et al.’s focus on GLBA notices by investigating how GLBA notices relate to other, more recently prevalent privacy notices (e.g., those required by CCPA). To our knowledge, this is the first study to comparatively examine the concise, two-page GLBA notice with the broader set of privacy policies provided by the same banks.
Studies examining CCPA and GDPR compliance (Tran et al., 2024a; O’Connor et al., 2021; Samarin et al., 2023) reveal that businesses frequently fail to implement legally required opt-out links and cookie controls. Even when these mechanisms are provided, manipulative design tactics are often used to discourage consumers from exercising their rights, in direct violation of CCPA’s restrictions (Tran et al., 2024b). Additionally, Aziz et al. (Aziz and Wilson, 2024) and Zimmeck et al. (Zimmeck et al., 2023) showed that automated privacy signals like Global Privacy Control (GPC), which are intended to provide consumers with an automated way to opt out, are frequently ignored by websites, despite legal mandates to honor them. Similarly, we found that some banks in our set ignored GPC signals even though they disclosed sale/sharing under the CCPA definition. Beyond interface-level manipulations, compliance gaps persist in actual data practices. Studies by Matte et al. (Matte et al., 2020) and Zhou et al. (Zhou et al., 2023) found that many apps and websites continue to collect personal data without proper disclosure, often using pre-selected consent options that subtly steer users toward agreement. The growing complexity of sector-specific and state-level privacy laws makes it increasingly difficult to maintain compliance and coherence. Our study examines the interplay between different compliance requirements through a consistency analysis, as consistency, clarity, and accessibility both within and across privacy documents provided by a single organization shape a consumer’s understanding and control of the use of their data.